OpenStudy (anonymous):

7x -7 =7 Find the solution of this equation,

OpenStudy (anonymous):

@ayegriffin

OpenStudy (anonymous):

Add 7 to both sides

OpenStudy (anonymous):

7x-7 + 7= 7+7 7 x =14 How much is x?

OpenStudy (anonymous):

easy its.......

OpenStudy (anonymous):

7x-7=7 +7 +7 7x=14 divide and get...